The energy levels of a particular quantum object are -11.8 eV, -5.5 eV, and -3.1 eV. If a collection of these objects is bombarded by an electron beam so that there are some objects in each excited state, what are the energies of the photons that will be emitted?
Answer #1

E2 -5.5 eV E3。. 3-1 ev mer of photons TA C mergさ 01 甲hntons eml(ied are
Answer #2

Solution: 

Given : E= - 11.8 eV ,  E= - 5.5  eV ,  E= - 3.1eV 

Energy of Photons

\DeltaE12 = E2 - E1 

=> \DeltaE12 = - 5.5 eV - (- 11.8 eV) 

=> \DeltaE12 = 6.3 eV 

 

\DeltaE13 = E3 - E1 

=> \DeltaE13 = - 3.1 eV - (- 11.8 eV) 

=> \DeltaE13 = 8.7 eV 

 

\DeltaE23 = E3 - E2 

=> \DeltaE23 = - 3.1eV - (- 5.5 eV) 

=> \DeltaE23 = 2.4 eV 

 

Energies of photon emitted will be 6.3 eV , 8.7 eV , 2.4 eV -------------Answer


answered by: Rajeev Kumar